# MSCloudCAM: Cross-Attention with Multi-Scale Context for Multispectral Cloud Segmentation

### v2

@@ Abstract
Clouds remain a critical challenge in optical satellite imagery, hindering reliable analysis for environmental monitoring, land cover mapping, and climate research. To overcome this, we propose MSCloudCAM, a Cross-Attention with Multi-Scale Context Network tailored for multispectral and multi-sensor cloud segmentation. Our framework exploits the spectral richness of Sentinel-2 (CloudSEN12) and Landsat-8 (L8Biome) data to classify four semantic categories: clear sky, thin cloud, thick cloud, and cloud shadow. MSCloudCAM combines a Swin Transformer backbone for hierarchical feature extraction with multi-scale context modules—ASPP and PSP—for enhanced scale-aware learning. A Cross-Attention block enables effective multisensor and multispectral feature fusion, while the integration of an Efficient Channel Attention Block (ECAB) and a Spatial Attention Module adaptively refine feature representations. Comprehensive experiments on CloudSEN12 and L8Biome demonstrate that MSCloudCAM delivers state-of-the-art segmentation accuracy, surpassing leading baseline architectures while maintaining competitive parameter efficiency and FLOPs. These results underscore the model’s effectiveness and practicality, making it well-suited for large-scale Earth observation tasks and real-world applications.
@@ I Introduction
Current deep learning–based cloud masking and segmentation methods have demonstrated strong performance by leveraging hierarchical feature representations. However, most existing approaches remain limited in their ability to generalize across different sensors and spectral configurations. This limitation becomes especially critical in multi-class scenarios that require discriminating between thin clouds, thick clouds, and cloud shadows, where subtle spectral and spatial variations are often misclassified.
Figure 1: High-level architecture of MSCloudCAM. The model integrates a Swin Transformer backbone with ASPP and PSP for multi-scale context, and incorporates a Cross Attention module for robust feature fusion across Sentinel-2 and Landsat-8 imagery.
We propose MSCloudCAM, a novel architecture designed for robust cloud segmentation in multispectral, multi-sensor imagery. The model exploits multi-scale context encoding and cross attention to effectively fuse complementary features from different resolutions and spectral domains, enabling accurate pixel-wise classification across heterogeneous datasets.
@@
  • We develop a hybrid backbone that integrates the Swin Transformer [1] with ASPP [2] and PSP [3] modules, enabling comprehensive multi-scale context aggregation.
  • We introduce a Cross Attention module that enhances feature fusion between high-level semantic and mid-level spatial features, improving robustness in complex cloud scenarios.
  • We demonstrate that MSCloudCAM achieves superior performance on both Sentinel-2 (CloudSEN12) [4] and Landsat-8 (L8Biome) [5] imagery, consistently outperforming strong state-of-the-art segmentation networks while maintaining computational efficiency.
@@ II-A Cloud Segmentation in Remote Sensing
Cloud segmentation in optical satellite imagery has long been studied due to its importance in atmospheric correction, land cover mapping, and environmental monitoring. Early rule-based methods, such as Fmask [6], applied band-specific thresholds and decision trees (e.g., cirrus, TIR bands in Landsat) but often misclassified mixed pixels, thin clouds, and bright surfaces like snow or sand [7].
Spectral index–based approaches leveraged indices such as NDSI, NDVI, and cloud probability measures [8], offering more robustness than simple thresholds but still failing under high spectral variability or semi-transparent clouds.
Later, machine learning classifiers (SVMs, Random Forests, gradient boosting) [9] combined spectral and textural features, improving accuracy but remaining limited by handcrafted inputs and shallow context modeling. These shortcomings motivated the transition to deep learning, which directly learns multi-scale, discriminative features from raw multispectral data.
@@ II-B Deep Learning Approaches
Deep learning has become the dominant paradigm for cloud segmentation in remote sensing, offering substantial improvements in accuracy, robustness, and generalization compared to traditional threshold-based methods.
CNN-based architectures remain foundational in this domain due to their strong spatial feature extraction capabilities and hierarchical representation learning. Encoder–decoder designs such as UNet [10] and its lightweight adaptation UNetMobv2 [11] preserve fine-grained localization via skip connections, making them effective for delineating cloud boundaries. DeepLabV3+ [2] enhances context capture through atrous convolutions and the Atrous Spatial Pyramid Pooling (ASPP) module, enabling multi-scale feature aggregation for detecting clouds of varying thickness. Several CNN variants have been developed specifically for cloud detection, including CDNetv1 [12] and CDNetv2 [13] with spectral index integration and multi-branch architectures, KappaMask [14] incorporating Kappa coefficient–based loss for class imbalance handling, DBNet [15] using dual-branch feature extraction for spatial–spectral fusion, SCNN [16] employing spectral–spatial separable convolutions, MCDNet [17] with multi-channel dense connections, and HRCloudNet [18] designed for high-resolution imagery.
Transformer-based architectures leverage self-attention mechanisms to capture long-range dependencies, a property highly beneficial for clouds that span large spatial extents or exhibit non-local spectral correlations. SegFormer [19] combines the lightweight MixVision Transformer (MiT) backbone with an MLP decoder for efficient global context modeling. Mask2Former [20] unifies instance, semantic, and panoptic segmentation through masked attention, offering strong adaptability to heterogeneous cloud patterns. UPerNet-InternImage [21] integrates a hierarchical transformer backbone (InternImage) with pyramid pooling for multi-scale representation, achieving state-of-the-art results on multiple remote sensing segmentation benchmarks.
Hybrid architectures aim to combine the strengths of CNNs and transformers by using CNN layers for low-level texture and spectral edge extraction, followed by transformer modules for global reasoning. Examples include HRNet–Transformer hybrids and cross-attention fusion models [1], which have been applied to atmospheric feature segmentation.
Foundation model–based segmentation has recently emerged as a promising direction, adapting large-scale pretrained vision transformers such as DINOv2 [22] for remote sensing segmentation. These models are trained on massive, diverse datasets and provide rich, generalizable features. While they have demonstrated strong zero-shot and fine-tuned performance in RGB domains, their adaptation to multispectral or multi-sensor data requires spectral-specific preprocessing and fine-tuning strategies.
Despite these advances, most existing models are trained and evaluated on single-sensor datasets, limiting their ability to generalize across multi-sensor imagery. Moreover, few explicitly integrate multi-scale spectral–spatial features with cross-attention mechanisms tailored for cloud segmentation. Our proposed MSCloudCAM addresses both gaps by combining a Swin Transformer[1] backbone with ASPP[2] and PSP[3] modules for multi-scale context aggregation, a Cross Attention block for sensor fusion, and ECAB[23] and spatial attention[24, 25] for adaptive feature refinement.
@@ II-C Gaps in Current Methods
Despite progress, most existing networks are trained on single-sensor datasets, limiting generalization to multi-sensor scenarios. Additionally, few works explicitly integrate multi-scale spectral-spatial features with cross-attention mechanisms for cloud segmentation. Our MSCloudCAM addresses both these gaps.
@@ III-A Overview
MSCloudCAM is a multispectral semantic segmentation network that integrates a Swin Transformer encoder with dual context modules, a cross-attention fusion mechanism, and a lightweight combined attention refinement block. The model is designed to capture long-range spectral–spatial dependencies while retaining fine-grained local structure. A multi-stage decoder with auxiliary supervision produces the final dense predictions. The overall pipeline is illustrated in Fig. 2.
@@ III-B Backbone Feature Extraction
Let $\mathbf{X}\in\mathbb{R}^{B\times C_{\text{in}}\times H_{0}\times W_{0}}$ denote the input multispectral image with $C_{\text{in}}$ spectral channels (13 for Sentinel-2, 11 for Landsat-8). The Swin Transformer encoder $\mathcal{E}(\cdot)$ produces a hierarchy of features:
@@ III-C Multi-Scale Context Modules
To enhance representational richness, we employ two complementary context encoders that operate at different feature hierarchies.
ASPP [2]
Applied to the deepest feature map $f_{4}$, Atrous Spatial Pyramid Pooling expands the effective receptive field through parallel dilated convolutions with multiple rates and an additional global context branch:
@@
PSP [3]
For the intermediate feature map $f_{3}$, Pyramid Scene Parsing aggregates multi-scale contextual cues by adaptive pooling at several grid scales:
@@ III-D Cross-Attention Fusion
The ASPP and PSP outputs are concatenated to form a joint query representation:
$x_{\text{cat}}=[\,x_{\text{ASPP}}\,\|\,x_{\text{PSP}}\,]\in\mathbb{R}^{B\times(C_{a}+C_{p})\times H\times W}.$ (4)
A convolutional multi-head cross-attention module then refines this fused tensor by treating $x_{\text{PSP}}$ as contextual guidance:
$Q$ $=W_{Q}*x_{\text{cat}},$ (5)
$K$ $=W_{K}*x_{\text{PSP}},$ (6)
$V$ $=W_{V}*x_{\text{PSP}},$ (7)
@@
$\tilde{x}$ $=W_{O}*\text{Attn}(Q,K,V)+x_{\text{cat}}.$ (9)
Here, * denotes convolution, $d_{k}$ is the per-head dimension, and residual addition ensures stable learning. This mechanism aligns global semantics from ASPP with fine-grained spatial cues from PSP.
@@ III-F Combined Attention Refinement
where $\phi_{\text{ECA}}$ denotes Efficient Channel Attention [23] and $\phi_{\text{SA}}$ is spatial attention [24, 25], and $\odot$ is element-wise multiplication. This highlights discriminative regions while preserving residual information.
@@ IV Experiments
TABLE I: Performance comparison on CloudSEN12 L1C (TOA).
@@ IV-A Datasets
CloudSEN12 [4] is a global Sentinel-2 cloud segmentation dataset with 13 spectral bands (10–60 m) and pixel-level labels for clear (0), thick cloud (1), thin cloud (2), and cloud shadow (3). We use the high subset, selecting 10k samples each from L1C and L2A. All bands are normalized by dividing top-of-atmosphere reflectance by $3000.0$.
@@ IV-D Evaluation Metrics
Per-class metrics ($\mathrm{IoU}_{i}$, $\mathrm{F1}_{i}$, $\mathrm{Acc}_{i}$) evaluate individual categories, while their means (mIoU, mF1, mAcc) are obtained by averaging across $K$ classes, giving equal weight regardless of class size. In contrast, overall accuracy (aAcc) is computed over all pixels, thus weighting larger classes more heavily.
Ignored labels.
Pixels with sentinel values (e.g., Fill in L8Biome) are removed from $\mathbf{C}$ and all denominators via an ignore_index mask, ensuring metrics are based on identical valid-pixel sets.
Reporting.
We report class-wise IoU and F1 (clear, thick cloud, thin cloud, cloud shadow), together with mIoU, mDice, mAcc, and aAcc. Unless noted, averages are macro over classes.
@@ V-A Quantitative Performance
We first evaluate the models quantitatively in terms of IoU, F1, and Accuracy on three benchmark datasets (CloudSEN12 L1C, CloudSEN12 L2A, and L8Biome). Tables IIV present the detailed results.
TABLE III: Performance comparison on CloudSEN12 L2A (BOA).
@@ V-B Qualitative Results
To further illustrate the effectiveness of the proposed method, we present visual comparisons of cloud segmentation results in Fig. 3. Each row corresponds to a different dataset, and each column corresponds to a different segmentation method. As shown in the figure, the proposed method produces more accurate delineation of thin clouds and cloud shadows, while reducing false detections compared to the baseline approaches.
@@
V-C Model Complexity Analysis
Table V summarizes the computational complexity of all evaluated models in terms of FLOPs and parameter counts. While some architectures (e.g., UPerNet-InternImage and Mask2Former-DINOv2) exhibit significantly higher complexity, our proposed MSCloudCAM achieves a favorable trade-off between accuracy and efficiency. Specifically, MSCloudCAM attains competitive performance while requiring fewer FLOPs than SegFormer-MiT-B5, and substantially fewer FLOPs than UPerNet-InternImage, with a moderate parameter count.
TABLE V: Model complexity comparison in terms of FLOPs and parameters.
Model FLOPs (G) Parameters (M)
U-Net-MobileNetV2 13.613 6.629
U-Net++-MobileNetV2 18.023 6.825
DeepLabV3+-MobileNetV2 6.057 4.379
U-Net-ResNet101 62.258 51.514
U-Net++-ResNet101 250.000 67.978
DeepLabV3+-ResNet101 56.192 45.670
UPerNet-InternImage-B-DCNv4 200–250 100–130
Mask2Former-Swin-T 73.090 47.400
Mask2Former-DINOv2-ViT-B/14 75–95 87–92
SegFormer-MiT-B5 55.430 45.652
MSCloudCAM 38.98 47.44
@@ VII Conclusion
We presented MSCloudCAM, a cross attention with multi-scale context network for multispectral, multi-sensor cloud segmentation. By integrating transformer-based backbones with multi-scale context and attention mechanisms, our method achieves superior segmentation accuracy over strong baselines on CloudSEN12 and L8Biome datasets. Future work will explore lightweight variants for onboard satellite processing and extend the model to spatiotemporal cloud tracking.

### v3

# MSCloudCAM: Multi-Scale Context Adaptation with Convolutional Cross-Attention for Multispectral Cloud Segmentation

@@ Abstract
Clouds remain a major obstacle in optical satellite imaging, limiting accurate environmental and climate analysis. To address the strong spectral variability and the large scale differences among cloud types, we propose MSCloudCAM, a novel multi-scale context adapter network with convolution based cross-attention tailored for multispectral and multi-sensor cloud segmentation. A key contribution of MSCloudCAM is the explicit modeling of multiple complementary multi-scale context extractors. And also, rather than simply stacking or concatenating their outputs, our formulation uses one extractor’s fine-resolution features and the other extractor’s global contextual representations enabling dynamic, scale-aware feature selection. Building on this idea, we design a new convolution-based cross attention adapter that effectively fuses localized, detailed information with broader multi-scale context. Integrated with a hierarchical vision backbone and refined through channel and spatial attention mechanisms, MSCloudCAM achieves strong spectral–spatial discrimination. Experiments on various multisensor datatsets e.g. CloudSEN12 (Sentinel-2) and L8Biome (Landsat-8) show that MSCloudCAM outperforms recent state-of-the-art models while maintaining competitive model complexity, highlighting the novelty and effectiveness of the proposed design for large-scale Earth observation.
@@ I Introduction
Recent deep learning–based cloud masking and segmentation methods have shown strong performance by leveraging hierarchical feature representations. However, most existing approaches struggle to generalize across heterogeneous sensors and spectral configurations. This limitation becomes especially critical in multi-class cloud segmentation requires modeling subtle spectral differences and large variations in spatial scale.
Figure 1: High-level architecture of MSCloudCAM. The model integrates a hierarchical vision backbone with dual multi-scale context extractors and introduces a novel cross-attention module.
To address these challenges, we propose MSCloudCAM, a cross-attention with multi-scale context network designed for robust cloud segmentation in multispectral, multi-sensor imagery. MSCloudCAM leverages complementary multi-scale context encoders and a newly designed cross-attention mechanism to dynamically fuse fine-grained spatial features with global contextual information. This design enables accurate pixel-wise classification across datasets with differing spectral responses, resolutions, and cloud characteristics.
@@
  • We introduce a hybrid multi-scale context framework that leverages two complementary multi-scale context extractors. One emphasizes on fine-grained spatial details and the other on capturing broad global semantics. Thus it enables us to achieve comprehensive multi-scale feature utilization capability from a hierarchical vision backbone.
  • We design a novel convolution-based cross-attention module, i.e. CAM that leverages these complementary feature hierarchies through an adaptive formulation. This mechanism enables dynamic, scale-aware feature fusion, offering a more principled alternative to conventional stacking or concatenation.
  • We demonstrate that MSCloudCAM achieves superior performance on both Sentinel-2 (CloudSEN12) [1] and Landsat-8 (L8Biome) [2], consistently outperforming state-of-the-art models while maintaining competitive computational efficiency.
@@ II-A Cloud Segmentation in Remote Sensing
Cloud segmentation is essential for atmospheric correction, land cover mapping, and environmental monitoring. Early rule-based methods such as Fmask [3] relied on band thresholds and decision trees but frequently misclassified thin clouds, mixed pixels, and bright surfaces. Spectral index–based approaches (e.g., NDSI, NDVI, cloud probability) [4] provided improved robustness yet still struggled under strong spectral variability or semi-transparent clouds. Traditional machine learning classifiers (SVM, Random Forests, boosting) [5] incorporated spectral–textural cues but remained constrained by handcrafted features and limited context modeling, motivating the shift toward deep learning.
@@ II-B Deep Learning Approaches
Deep learning has become the dominant strategy for cloud segmentation due to its ability to model spatial–spectral structure directly from multispectral inputs.
CNN-based architectures such as UNet [6] and UNetMobv2 [7] capture fine localization through skip connections.DeepLabV3+ [8] enhances multi-scale context via dilated convolutions and ASPP. Numerous CNN variants target cloud detection specifically, including CDNetv1/v2 [9, 10], KappaMask [11], DBNet [12], SCNN [13], MCDNet [14], and HRCloudNet [15], each improving spectral–spatial fusion or multi-scale representation.
Transformer-based methods leverage self-attention for long-range dependency modeling, crucial for large or spatially diffuse clouds. SegFormer [16] uses a lightweight MiT backbone for global context, while Mask2Former [17] delivers strong generalization through masked attention. UPerNet-InternImage [18] combines hierarchical transformers with pyramid pooling to achieve state-of-the-art results in remote sensing segmentation.
Hybrid models merge CNNs for low-level edge and texture extraction with transformer blocks for global reasoning, including HRNet–Transformer variants and cross-attention designs [19].
Foundation-model–based segmentation adapts pretrained vision transformers such as DINOv2 [20] to remote sensing; however, effective use of multispectral data still requires spectral alignment and domain-specific fine-tuning.
@@ II-C Gaps in Current Methods
Despite significant progress, most approaches are designed for single-sensor datasets, limiting cross-sensor generalization. Moreover, few models explicitly integrate complementary multi-scale spectral–spatial cues using directional cross-attention. Our MSCloudCAM addresses both challenges.
@@ III-A Overview
MSCloudCAM is a multispectral semantic segmentation network that integrates a hierarchical vision backbone e.g. Swin Transformer [19] encoder with dual context extractor modules, a cross-attention fusion mechanism, and a lightweight combined attention refinement block. The model is designed to capture long-range spectral–spatial dependencies while retaining fine-grained local structure. A multi-stage decoder with auxiliary supervision produces the final dense predictions. The overall pipeline is illustrated in Fig. 2.
@@ III-B Backbone Feature Extraction
Let $\mathbf{X}\in\mathbb{R}^{B\times C_{\text{in}}\times H_{0}\times W_{0}}$ denote the input multispectral image with $C_{\text{in}}$ spectral channels (13 for Sentinel-2, 11 for Landsat-8). The encoder $\mathcal{E}(\cdot)$ produces a hierarchy of features:
@@ III-C Multi-Scale Context Modules
To enhance representational richness, we employ two complementary multi context encoders that operate at different feature hierarchies.
As a part of dual multiscale context extractor, ASPP [8] is applied to the deepest feature map $f_{4}$, Atrous Spatial Pyramid Pooling expands the effective receptive field through parallel dilated convolutions with multiple rates and an additional global context branch:
@@
For the intermediate feature map $f_{3}$, Pyramid Scene Parsing i.e. PSP [21] aggregates multi-scale contextual cues by adaptive pooling at several grid scales:
@@ III-D Cross-Attention Fusion
The outputs from the two multi-scale context extractors are concatenated to form a joint feature representation:
$x_{\text{cat}}=[\,x_{\text{MC1}}\,\|\,x_{\text{MC2}}\,]\in\mathbb{R}^{B\times(C_{a}+C_{p})\times H\times W},$ (4)
where $x_{\text{MC1}}$ emphasizes fine spatial detail from ASPP module and $x_{\text{MC2}}$ from PSP module encodes broader contextual semantics. To effectively combine these complementary cues, we employ a convolutional cross-attention mechanism. Here, $x_{\text{cat}}$ is used to generate the queries, as the concatenated representation benefits from selectively incorporating global context to refine local structure. Meanwhile, $x_{\text{MC2}}$ provides the keys and values, since its pooled multi-scale features offer stable, region-level guidance suitable for modulating finer details. This directional interaction allows each spatial location in $x_{\text{MC1}}$ to attend to the most relevant semantic context encoded by $x_{\text{MC2}}$.
Formally, the attention is computed as follows: the query $Q=W_{Q}*x_{\text{cat}}$, key $K=W_{K}*x_{\text{MC2}}$, and value $V=W_{V}*x_{\text{MC2}}$ are first obtained.
@@
$\tilde{x}$ $=W_{O}*\text{Attn}(Q,K,V)+x_{\text{cat}},$ (6)
where * denotes convolution and residual addition preserves the original spatial representation. This formulation enables fine-grained features to be adaptively enhanced by global contextual cues, producing a richer and more discriminative multi-scale representation.
@@ III-F Combined Attention Refinement
where $\phi_{\text{ECA}}$ denotes Efficient Channel Attention [22] and $\phi_{\text{SA}}$ denotes spatial attention [23, 24]. This joint formulation first reweights channels based on their global importance and then emphasizes spatially salient regions, allowing the model to focus on both discriminative spectral responses and structurally meaningful areas.
@@ IV Experiments
TABLE I: Performance comparison on CloudSEN12 L2A (BOA).
@@ IV-A Datasets
CloudSEN12 [1] is a global Sentinel-2 cloud segmentation dataset with 13 spectral bands (10–60 m) and pixel-level labels for clear (0), thick cloud (1), thin cloud (2), and cloud shadow (3). We use the high subset, selecting 10k samples each from L1C and L2A. L1C provides Top-of-Atmosphere (TOA) reflectance, while L2A provides Bottom-of-Atmosphere (BOA) surface reflectance. All bands are normalized by dividing top-of-atmosphere reflectance by $3000.0$.
@@ IV-D Evaluation Metrics
Class-wise metrics ($\mathrm{IoU}_{i}$, $\mathrm{F1}_{i}$, $\mathrm{Acc}_{i}$) evaluate individual categories, while their means (mIoU, mF1, mAcc) are obtained by averaging across $K$ classes, giving equal weight regardless of class size. In contrast, overall accuracy (aAcc) is computed over all pixels, thus weighting larger classes more heavily.
@@ V-A Quantitative Performance
We first evaluate the models quantitatively in terms of IoU, F1, and Accuracy on three benchmark datasets (CloudSEN12 L1C, CloudSEN12 L2A, and L8Biome). Tables IIIIV present the detailed results.
TABLE III: Performance comparison on CloudSEN12 L1C (TOA).
@@ V-B Qualitative Results
To further demonstrate the effectiveness of our method, Fig. 3 presents visual comparisons of cloud segmentation results. Each row shows a different dataset, and each column a different method.
@@
V-C Ablation Study
To evaluate the role of each multicontext component, we conduct ablation experiments by selectively replacing or removing modules. Replacing ASPP with PSP enhances global context but loses fine detail, while using ASPP instead of PSP improves multi-scale local features but sacrifices global scene understanding.
Using PSP and ASPP together provides complementary global–local context and outperforms either module alone. Removing the Cross-Attention Module and using simple concatenation weakens the model’s ability to emphasize informative regions. The full MSCloudCAM system with PSP, ASPP, and CAM delivers the highest accuracy.
TABLE V: Ablation study evaluating dual multi context extractor configurations and our cross-attention module (CAM) on CloudSEN12 L2A (BOA).
Dual Extractor Fusion Type mIoU mDice mAcc aAcc
PSP + PSP CAM 70.26 81.61 92.05 86.41
ASPP + ASPP CAM 70.09 81.25 91.93 86.18
PSP + ASPP Concat 71.48 82.44 93.52 88.02
PSP + ASPP CAM 73.12 83.70 94.85 89.69
V-D Model Complexity Analysis
Table VI summarizes the computational complexity of all evaluated models in terms of FLOPs, parameter counts, and inference efficiency. In addition to latency (ms/image), we report throughput (TP, samples/s) to quantify batch inference performance. All experiments were conducted on a single NVIDIA A100 40GB GPU using PyTorch 2.3 and CUDA 12.1, with a batch size of 8 and an input resolution of 512×512. All models were evaluated in FP32 precision. Our proposed MSCloudCAM achieves a noteworthy balance between accuracy and computational efficiency and CAM module only adds 1.7M trainable parameters to the model.
TABLE VI: Model complexity comparison in terms of FLOPs (G), parameters (M), inference time (ms/image), and TP as throughput (FPS).
[…]
U-Net-MobileNetV2 13.6 6.6 11.5 87.0
DeepLabV3+-MobileNetV2 6.1 4.3 14.8 67.5
DeepLabV3+-ResNet101 56.1 45.6 38.5 26.0
UPerNet-InternImageDCNv4 193.5 110.8 52.0 19.2
Mask2Former-Swin-T 73.09 47.4 33.6 29.7
Mask2Former-DINOv2-B14 85.6 90.6 44.0 22.7
SegFormer-MiT-B5 55.4 45.6 28.7 34.8
MSCloudCAM 38.98 47.44 15.9 62.7
@@ VII Conclusion
We presented MSCloudCAM, a cross attention with multi-scale context network for multispectral, multi-sensor cloud segmentation. By integrating hierarchical vision backbone with dual multi-scale context extractors and emlpoying our novel convolution based cross attention module, our method achieves superior segmentation accuracy over state-of-art-models on CloudSEN12 and L8Biome datasets. Future work will explore lightweight variants for onboard satellite processing and extend the model to spatiotemporal cloud tracking.